OpenStudy (anonymous):

When adding and subtracting radicals, can you join like terms ignoring the signs

OpenStudy (amistre64):

??

OpenStudy (anonymous):

what do u mean by signs?

OpenStudy (angela210793):

what r radicals? O.o

OpenStudy (anonymous):

\[2\sqrt{10}-3\sqrt{40}\]

OpenStudy (amistre64):

when adding and subtracting radicals; the radicals act like a variable in that you can only work them out if they are the same...

OpenStudy (amistre64):

you have to rewrite sqrt(40) in terms of sqrt(10)s to work this out

OpenStudy (amistre64):

sqrt(40) = sqrt(10*4) = sqrt(10) sqrt(4) sqrt(10) [2 - 3(2)] = -4sqrt(10)

OpenStudy (angela210793):

2sqr10-3sqr(4*10)=2sqr10-6sqr10=-4sqr10.....i guess :)